OpenStudy (anonymous):

Create your own two-step equation with NO SOLUTION as the solution. Show the steps to solving the equation and explain in words why your work proves that there are NO SOLUTIONS.

OpenStudy (inkyvoyd):

y=ax+b y=ax+c Where a, b, and c are constants, and b=/=c

OpenStudy (inkyvoyd):

Then, rearrange the two equations so they look unalike.

OpenStudy (anonymous):

still not getting this :/

OpenStudy (inkyvoyd):

Sorry, I read the problem wrong.

OpenStudy (anonymous):

oh okay lol i was sooo confused

OpenStudy (inkyvoyd):

make an equation of the form ax+b=ax+c b=/=c There is no solution to this equation because the ax's cancel out.

OpenStudy (anonymous):

can you rewrite that

OpenStudy (inkyvoyd):

Basically we want to make an equation with no solution as the result. In other words, we wind up with two constants, (let's say b and c), and these constants are not equal. Start with b=c (not true, but we are working backwards) ax+b=ax+c (still not true, but this is what the question was asking)

OpenStudy (anonymous):

ohhh so all i need to do is make a contradiction!

OpenStudy (inkyvoyd):

in a way, yes. Let me give you an example. If we have 3x+5=3x-2 It looks at first that we can solve the equation. But then we cancel the x's out 5=-2 This is obviously wrong